First Perfect Square

Given two integers (M and N), write a program to print the first perfect square in a given range.Input


The first line of input will contain a positive integer (M).

The second line of input will contain a positive integer (N).Output


The output should be a single line containing the first perfect square in a given range.Explanation


For example, if the given numbers are M is 4 and N is 16, the perfect squares from 4 to 16 are 4(2 * 2), 9(3 * 3), and 16(4 * 4), as the first perfect square is 4. So the output should be 4.


Sample Input 1

4

16


Sample Output 1

4


Sample Input 2

5

8


Sample Output 2

No Perfect Square




1
Expert's answer
2021-03-27T00:16:53-0400
x = int(input("First number:"))
y = int(input("Second number:"))
flag = False
for i in range (x, y + 1):
    if not flag:
        j = 1;
        while j * j <= i:
            if j * j == i:
                 print(j*j)
                 flag = True
                 break
            j = j + 1
        i = i + 1
    else:
        break
    
if not flag:
    print("No perfect square")